A certain cross of rat will result in white-black in the ratio of 8:5.?
a) What is the probability that among 9 such offspring, 5 will be white? (accurate to 4 decimal places)
b) What is the expected number of black rats from part a)?

Answer by Boreal(15235)

probability of white is (8/13)
assuming independence, this is 9C5 (8/13)^5*(5/13)^4. Use fractions for the exact value.
=0.2433.
-
This is Np=9*5/13=45/13, or about 3.5